Child's Play Company makes a plastic rattle for toddlers. The rattle is generally marketed through exclusive retailers located in upscale shopping malls. The company used 23 retailers. In late 2016, Diana Suares, the president of the company, was considering alternative marketing plans presented to her by Bill Dings, the marketing manager. Based on sales from January through October 2016, Diana expected that 2016 sales would amount to 300,000 units at $8.00 per unit. Diana Suares also had with her some cost data for 2016 supplied by the CFO, Don Capp. Don expects that these costs are reliable estimates for a volume up to 400,000 units. Beyond 400,000 units, the company would have to rent additional machines (with a capacity of 100,000 units each) at an annual cost of $50,000 per machine (in 2016). The data are presented in Exhibit A. Total fixed manufacturing and selling and administrative costs are each expected to increase in 2017 by 10 percent because of inflation. Variable costs per unit and the selling price would stay the same as in 2016.

Exhibit A (2016 Cost Data) Manufacturing costs for rattles (based on production volume of 300,000 units).

Direct material $0.80 per unit *

Direct labor $10 per hour *

(Each worker can make 20 units in an hour)

Packaging $0.75 per unit *

Power, supplies, indirect labor, and other variable production costs $1.20 per unit *

Supervisory salaries, equipment rental, and miscellaneous production costs $1.80 per unit *

These costs vary in total with production volume

Selling and administrative costs (based on sales volume of 300,000 units).

Commissions to sales staff 10% of price to retailer **

Shipping $0.50 per unit **

Advertising and promotion $0.60 per unit

Administrative staff salaries, depreciationon office equipment, etc $0.90 per unit

** These costs vary in total with sales volume.

10. What is the contribution margin per unit?

11. What is the contribution margin if 1000 units are sold?

12. What is the contribution margin if 100,000 units are sold?

13. What is the profit at the break even volume?

14. What is the contribution margin at the break even volume?

15. What is the break even volume?

16. What is the profit at 1000 units above break even?

17. What is the profit at 2000 units below the break even?

18. What is the profit at 100,000 units above break even?

19. What is the margin of safety at 1000 units above break even?

20. What is the sales volume required to give $100,000 profit before tax?

21. What is the sales volume required to give $180,000 after tax, assuming a tax rate of 40%?
Answer #1
Supervisory salaries, equipment rental, and miscellaneous production costs $       594,000 =300000*1.8*1.1
Advertising and promotion $       198,000 =300000*0.6*1.1
Administrative staff salaries, depreciation office equipment $       297,000 =300000*0.9*1.1
Total Fixed Costs $   1,089,000
Direct Material $       240,000 =300000*0.8
Direct Labor $       150,000 =300000*10/20
Packaging $       225,000 =300000*0.75
Power, supplies, indirect labor, and other variable production costs $       360,000 =300000*1.2
Commissions to sales $       240,000 =300000*8*10%
Shipping $       150,000 =300000*0.5
Total Variable Costs $   1,365,000


Variable Cost per unit = $1365000 / 300000 = $4.55 per unit

10. Contribution margin per unit = Sales Price - Variable cost per unit
= $8 - 4.55 = $3.45 per unit

11. Contribution margin for 1000 units = $3.45 x 1000 = $3450

12. Contribution margin for 100000 units = $3.45 x 100000 = $345000

13. Profit at break even volume = $0
Break even itself means when revenues is equal to expenses

14. Contribution margin at breakeven volume = fixed costs = $1089000

15. Break even volume = Fixed Costs / Contribution Margin per unit
= $1089000 / 3.45 = 315653 units

As per HOMEWORKLIB RULES we are supposed to answer 4 parts, i have answered 6. so kindly post other parts separately